# Intrinsic Training Signals for Federated Learning Aggregation

### v1

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy. While existing approaches for aggregating client-specific classification heads and adapted backbone parameters require architectural modifications or loss function changes, our method uniquely leverages intrinsic training signals already available during standard optimization. We present LIVAR (Layer Importance and VARiance-based merging), which introduces: i) a variance-weighted classifier aggregation scheme using naturally emergent feature statistics, and ii) an explainability-driven LoRA merging technique based on SHAP analysis of existing update parameter patterns. Without any architectural overhead, LIVAR achieves state-of-the-art performance on multiple benchmarks while maintaining seamless integration with existing FL methods. This work demonstrates that effective model merging can be achieved solely through existing training signals, establishing a new paradigm for efficient federated model aggregation. The code will be made publicly available upon acceptance.
@@ 3 Preliminaries
[…]
@@ 4 LIVAR
For the backbone, we employ Low-Rank Adaptation with our novel Proxy-Calibrated Layer Importance Weighting mechanism, using explainability-driven coefficients to merge client updates. The classification head uses data-variance-aware merging that automatically adapts to client data heterogeneity, ensuring robustness under non-IID distributions. The dual merging strategy is depicted in Figure 1.
@@ 4.1 Proxy-Calibrated Layer Importance Weighting
Third, we employ SHAP analysis [15] to interpret the trained model (see Figure 2), revealing distinct roles of the $A$ and $B$ matrices – the former initialized via Kaiming uniform distribution [4] to explore directions, the latter initialized to zeros to control direction utilization. The SHAP values are discretized through a quantile-based function $G_{SHAP}$ to yield the final coefficients:
@@
Additional implementation details regarding the quantile-based discretization procedure are provided in Appendix 0.A.
@@
     $\Delta W_{m}\leftarrow(training)$triangleright  Collect $(\omega_{m}^{l_{A}},\omega_{m}^{l_{B}})$ during training
@@
     $\alpha_{m}^{l}\leftarrow$ use Equation 5
     $\Delta W^{l}\leftarrow$ use Equation 3 $\triangleright$  Proxy-Calibrated Layer Importance Weighting
@@
     $H^{(c)}\leftarrow$ use Equation 6 $\triangleright$  Head Class-Specific Variance Weighting
@@
Beyond the proxy dataset. The function $G_{SHAP}(\cdot)$, which estimates merging coefficients from importance signals $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ observed during training, enables federated aggregation across arbitrary datasets. Our quantile-based discretization ensures robust generalization beyond the proxy dataset used for calibration. We validate this generalization capability through a posteriori analysis of SHAP value distributions across multiple benchmarks in Section 5.4.
For server-side aggregation, the coefficients $\alpha_{m}^{l}$ computed via Equation 5 are applied to combine client updates $\Delta W_{m}^{l}$, as in Equation 3, with each client’s importance measures $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ derived from their local training process.
@@ 4.2 Head Class-specific Variance Weighting
This variance-weighted scheme prioritizes classification heads whose training data exhibits broader coverage of each class’s feature distribution. The complete training procedure is formally specified in Algorithm 1.
@@ 5.2 Results of LIVAR
In distribution. As reported in Table 1, for both in-distribution datasets, we observe consistent trends: LIVAR achieves state-of-the-art results across all $\beta$ settings, while CCVR consistently ranks second. For both datasets, higher degrees of distribution imbalance yield more pronounced differences in performance – while LIVAR and CCVR maintain their leading performance, the model merging approaches follow closely, and other methods fail to remain robust under these harder conditions.
@@
Out of distribution. Table 2 shows that the EuroSAT dataset proves less informative for evaluating Federated Learning with ViT-based methods, as all models achieve similar accuracy levels – likely due to the dataset’s limited $10$-class structure. In contrast, CUB-200 presents a more challenging benchmark where LIVAR and CCVR significantly outperform competing approaches. LIVAR achieves state-of-the-art results for moderate data imbalance ($\beta\in\{1,0.5\}$), with CCVR consistently ranking second. However, under extreme imbalance ($\beta=0.2$), CCVR surpasses LIVAR, demonstrating the advantages of its classifier rebalancing strategy over our variance-based approach.
Method composability. We investigate LIVAR’s compatibility with existing federated learning techniques through integration with CCVR on CUB-200 – a dataset representing a significant domain shift from ViT’s pre-training distribution. While CCVR alone outperforms LIVAR in the extreme imbalance setting ($\beta=0.2$), we demonstrate their complementary strengths when combined: applying LIVAR’s layer-specific aggregation followed by CCVR’s server-side calibration with synthetic data yields superior performance to either method in isolation. As shown in Table 4, this synergy highlights how LIVAR’s feature variance head weighting provides foundational benefits that enhance subsequent refinement techniques like CCVR. This composability represents a significant advantage, as it allows practitioners to combine multiple techniques to address the unique challenges of different federated learning scenarios.
@@ 5.3 Ablation Study on LIVAR Components
We systematically evaluate LIVAR’s components by measuring Average Accuracy and standard deviation on CIFAR-100 with extreme label imbalance ($\beta=0.05$). Starting with a baseline of FedAvg-aggregated ViT-B/16 parameters – including both LoRA residuals $\Delta W^{l}$ and the classification head $H$Table 4 demonstrates progressive improvements through incremental enhancements. The Proxy-Calibrated Layer Importance Weighting ($\alpha^{l}$) achieves statistically significant accuracy gains beyond the standard deviation across five runs (reported in parentheses in Table 4). Subsequent incorporation of Head Class-Specific Variance Weighting ($\sigma^{(c)}$) delivers substantially greater improvement, corroborating the classification head’s critical role. The full LIVAR framework combines these components synergistically to achieve peak performance.
@@ 5.4 A Posteriori Analysis of SHAP Results
To verify the generalizability of our proxy-calibrated SHAP analysis explained in Section 4.1, we repeated the calibration procedure across all four datasets: CIFAR-100, ImageNet-R, EuroSAT, and CUB-200. Importantly, this analysis does not involve recalibrating the $G_{SHAP}(\cdot)$ function for each specific dataset – since such per-dataset calibration would be impossible in a federated setting where the joint data distribution is inaccessible – but instead demonstrates that the original proxy-based calibration effectively generalizes to different data domains.
Following the same methodology that produced the results in Figure 2, we generated comparable analyses for each dataset. As shown in Figure 3 for ImageNet-R, the SHAP value trends remain consistent: $\omega^{l_{A}}$ exhibits decreasing or stable influence with increasing feature values, while $\omega^{l_{B}}$ maintains its clear increasing trend across all datasets. These consistent patterns (fully documented in Appendix 0.B) confirm the proxy dataset’s representativeness, despite variations in specific impact magnitudes.
@@ Appendix 0.A Details on $G_{SHAP}(\omega^{l_{A}},\omega^{l_{B}})$
At the conclusion of each round, the server calculates the quantile values $Q_{A}$ and $Q_{B}$ for the complete distribution of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$, respectively, across all clients $m$ and layers $l$. Subsequently, for each layer, the corresponding value is assigned based on the intervals specified in Table A, and a normalization is applied across each layer to ensure that the sum of the resulting SHAP values after the discretization of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ equals $1$. This procedure defines the function $G_{SHAP}(\cdot)$, which is executed server-side at the end of each communication round, returning $\alpha^{l}$ coefficients.
@@ Appendix 0.B Robustness in the Choice of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$
Specifically, we initially trained the Random Forest using three features: $\omega_{m}^{l_{A}}$, $\omega_{m}^{l_{B}}$, and the number of training samples for each client. When evaluating the SHAP values of these three features, it became evident that the impact of the number of samples is negligible, as shown in Figure A. This result is directly attributable to the fact that $\lambda_{m}^{l}$ coefficients are estimated layer-wise. More importantly, this finding supports our hypothesis that $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ are sufficient and appropriate choices for predicting optimal merging coefficients.
@@ Appendix 0.C Additional a posteriori Analysis of SHAP Results
As demonstrated in the main paper, the SHAP trend observed for the $\omega^{l_{A}}$ and $\omega^{l_{B}}$ quantiles is consistent between the Tiny-ImageNet proxy dataset and other distributions. Figures Fig. B, Fig. C, and Fig. D further illustrate that this trend is consistently maintained across the CIFAR-100, EuroSAT, and CUB-200 datasets. Specifically, for CIFAR-100, the negative impact of $\omega^{l_{A}}$ at high quantiles is more pronounced at the upper extreme of the distribution. The impact of $\omega^{l_{B}}$ is generally similar across the datasets examined, with the notable exception of CUB-200, where its relative impact is lower in magnitude. Consequently, the non-decreasing trend for $\omega^{l_{B}}$ is not fully maintained across the entire distribution for this dataset. This observation is likely attributable to the dissimilarity between the CUB-200 dataset and the pretraining data, suggesting that the change in impact layer direction, represented by the $A$ matrices, exerts a more significant influence on the optimal merging strategy in this case.

### v2

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy. While existing approaches for aggregating client-specific classification heads and adapted backbone parameters require architectural modifications or loss function changes, our method uniquely leverages intrinsic training signals already available during standard optimization. We present LIVAR (Layer Importance and VARiance-based merging), which introduces: i) a variance-weighted classifier aggregation scheme using naturally emergent feature statistics, and ii) an explainability-driven LoRA merging technique based on SHAP analysis of existing update parameter patterns. Without any architectural overhead, LIVAR achieves state-of-the-art performance on multiple benchmarks while maintaining seamless integration with existing FL methods. This work demonstrates that effective model merging can be achieved solely through existing training signals, establishing a new paradigm for efficient federated model aggregation. The code is available at https://github.com/aimagelab/fed-mammoth.
@@ 3 Preliminaries
$\underset{\theta}{\text{minimize}}\quad\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{(x,y)\sim\mathcal{D}_{m}}\mathcal{L}(f_{\theta}(x),y)$ (2)
@@ 4 LIVAR
For the backbone, we employ Low-Rank Adaptation with our novel Proxy-Calibrated Layer Importance Weighting mechanism, using explainability-driven coefficients to merge client updates. The classification head uses data-variance-aware merging that automatically adapts to client data heterogeneity, ensuring robustness under non-IID distributions. The dual merging strategy is depicted in Figure˜1.
@@ 4.1 Proxy-Calibrated Layer Importance Weighting
Third, we employ SHAP analysis [15] to interpret the trained model (see Figure˜2), revealing distinct roles of the $A$ and $B$ matrices – the former initialized via Kaiming uniform distribution [4] to explore directions, the latter initialized to zeros to control direction utilization. The SHAP values are discretized through a quantile-based function $G_{SHAP}$ to yield the final coefficients:
@@
Additional implementation details regarding the quantile-based discretization procedure are provided in Appendix˜0.A.
@@
  $\Delta W_{m}\leftarrow(training)$ $\triangleright$  Collect $(\omega_{m}^{l_{A}},\omega_{m}^{l_{B}})$ during training
@@
  $\alpha_{m}^{l}\leftarrow$ use Equation˜5
  $\Delta W^{l}\leftarrow$ use Equation˜3 $\triangleright$  Proxy-Calibrated Layer Importance Weighting
@@
  $H^{(c)}\leftarrow$ use Equation˜6 $\triangleright$  Head Class-Specific Variance Weighting
@@
Beyond the proxy dataset. The function $G_{SHAP}(\cdot)$, which estimates merging coefficients from importance signals $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ observed during training, enables federated aggregation across arbitrary datasets. Our quantile-based discretization ensures robust generalization beyond the proxy dataset used for calibration. We validate this generalization capability through a posteriori analysis of SHAP value distributions across multiple benchmarks in Section˜5.4.
For server-side aggregation, the coefficients $\alpha_{m}^{l}$ computed via Equation˜5 are applied to combine client updates $\Delta W_{m}^{l}$, as in Equation˜3, with each client’s importance measures $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ derived from their local training process.
@@ 4.2 Head Class-specific Variance Weighting
This variance-weighted scheme prioritizes classification heads whose training data exhibits broader coverage of each class’s feature distribution. The complete training procedure is formally specified in Algorithm˜1.
@@ 5.2 Results of LIVAR
In distribution. As reported in Table˜1, for both in-distribution datasets, we observe consistent trends: LIVAR achieves state-of-the-art results across all $\beta$ settings, while CCVR consistently ranks second. For both datasets, higher degrees of distribution imbalance yield more pronounced differences in performance – while LIVAR and CCVR maintain their leading performance, the model merging approaches follow closely, and other methods fail to remain robust under these harder conditions.
@@
Out of distribution. Table˜2 shows that the EuroSAT dataset proves less informative for evaluating Federated Learning with ViT-based methods, as all models achieve similar accuracy levels – likely due to the dataset’s limited $10$-class structure. In contrast, CUB-200 presents a more challenging benchmark where LIVAR and CCVR significantly outperform competing approaches. LIVAR achieves state-of-the-art results for moderate data imbalance ($\beta\in\{1,0.5\}$), with CCVR consistently ranking second. However, under extreme imbalance ($\beta=0.2$), CCVR surpasses LIVAR, demonstrating the advantages of its classifier rebalancing strategy over our variance-based approach.
Method composability. We investigate LIVAR’s compatibility with existing federated learning techniques through integration with CCVR on CUB-200 – a dataset representing a significant domain shift from ViT’s pre-training distribution. While CCVR alone outperforms LIVAR in the extreme imbalance setting ($\beta=0.2$), we demonstrate their complementary strengths when combined: applying LIVAR’s layer-specific aggregation followed by CCVR’s server-side calibration with synthetic data yields superior performance to either method in isolation. As shown in Table˜4, this synergy highlights how LIVAR’s feature variance head weighting provides foundational benefits that enhance subsequent refinement techniques like CCVR. This composability represents a significant advantage, as it allows practitioners to combine multiple techniques to address the unique challenges of different federated learning scenarios.
@@ 5.3 Ablation Study on LIVAR Components
We systematically evaluate LIVAR’s components by measuring Average Accuracy and standard deviation on CIFAR-100 with extreme label imbalance ($\beta=0.05$). Starting with a baseline of FedAvg-aggregated ViT-B/16 parameters – including both LoRA residuals $\Delta W^{l}$ and the classification head HHTable˜4 demonstrates progressive improvements through incremental enhancements. The Proxy-Calibrated Layer Importance Weighting ($\alpha^{l}$) achieves statistically significant accuracy gains beyond the standard deviation across five runs (reported in parentheses in Table˜4). Subsequent incorporation of Head Class-Specific Variance Weighting ($\sigma^{(c)}$) delivers substantially greater improvement, corroborating the classification head’s critical role. The full LIVAR framework combines these components synergistically to achieve peak performance.
@@ 5.4 A Posteriori Analysis of SHAP Results
To verify the generalizability of our proxy-calibrated SHAP analysis explained in Section˜4.1, we repeated the calibration procedure across all four datasets: CIFAR-100, ImageNet-R, EuroSAT, and CUB-200. Importantly, this analysis does not involve recalibrating the $G_{SHAP}(\cdot)$ function for each specific dataset – since such per-dataset calibration would be impossible in a federated setting where the joint data distribution is inaccessible – but instead demonstrates that the original proxy-based calibration effectively generalizes to different data domains.
Following the same methodology that produced the results in Figure˜2, we generated comparable analyses for each dataset. As shown in Figure˜3 for ImageNet-R, the SHAP value trends remain consistent: $\omega^{l_{A}}$ exhibits decreasing or stable influence with increasing feature values, while $\omega^{l_{B}}$ maintains its clear increasing trend across all datasets. These consistent patterns (fully documented in Appendix˜0.B) confirm the proxy dataset’s representativeness, despite variations in specific impact magnitudes.
@@ Appendix 0.A Details on $G_{SHAP}(\omega^{l_{A}},\omega^{l_{B}})$
At the conclusion of each round, the server calculates the quantile values $Q_{A}$ and $Q_{B}$ for the complete distribution of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$, respectively, across all clients $m$ and layers $l$. Subsequently, for each layer, the corresponding value is assigned based on the intervals specified in Table˜A, and a normalization is applied across each layer to ensure that the sum of the resulting SHAP values after the discretization of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ equals $1$. This procedure defines the function $G_{SHAP}(\cdot)$, which is executed server-side at the end of each communication round, returning $\alpha^{l}$ coefficients.
@@ Appendix 0.B Robustness in the Choice of $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$
Specifically, we initially trained the Random Forest using three features: $\omega_{m}^{l_{A}}$, $\omega_{m}^{l_{B}}$, and the number of training samples for each client. When evaluating the SHAP values of these three features, it became evident that the impact of the number of samples is negligible, as shown in Figure˜A. This result is directly attributable to the fact that $\lambda_{m}^{l}$ coefficients are estimated layer-wise. More importantly, this finding supports our hypothesis that $\omega_{m}^{l_{A}}$ and $\omega_{m}^{l_{B}}$ are sufficient and appropriate choices for predicting optimal merging coefficients.
@@ Appendix 0.C Additional a posteriori Analysis of SHAP Results
As demonstrated in the main paper, the SHAP trend observed for the $\omega^{l_{A}}$ and $\omega^{l_{B}}$ quantiles is consistent between the Tiny-ImageNet proxy dataset and other distributions. Figures Fig.˜B, Fig.˜C, and Fig.˜D further illustrate that this trend is consistently maintained across the CIFAR-100, EuroSAT, and CUB-200 datasets. Specifically, for CIFAR-100, the negative impact of $\omega^{l_{A}}$ at high quantiles is more pronounced at the upper extreme of the distribution. The impact of $\omega^{l_{B}}$ is generally similar across the datasets examined, with the notable exception of CUB-200, where its relative impact is lower in magnitude. Consequently, the non-decreasing trend for $\omega^{l_{B}}$ is not fully maintained across the entire distribution for this dataset. This observation is likely attributable to the dissimilarity between the CUB-200 dataset and the pretraining data, suggesting that the change in impact layer direction, represented by the $A$ matrices, exerts a more significant influence on the optimal merging strategy in this case.